OpenStudy (anonymous):

In 5 years, Leroy will be twice as old as his brother. Today, their ages add up to 14. How old is Leroy? Help?

OpenStudy (shadowfiend):

You can set this up as a system of equations. Let's take the two parts of this problem: In 5 years, Leroy will be twice as old as his brother. If we make Leroy's age \(x\) and his brother's age \(y\), this means: \[x + 5 = 2(y + 5)\] This says `Leroy's age plus five is the same as two times his brother's age plus five' -- meaning in 5 years, Leroy's age will be two times that of his brother's in five years. That's the first part of the system. The second part is that today, their ages add up to 14. This means: \[x + y = 14\] Can you solve that?

OpenStudy (anonymous):

its 3 and 11